# Revisiting CMSSM with Non-Universal Gaugino Masses under Current Constraints

### v1

@@ I Introduction
These phenomena are difficult to explain within the framework of the SM and strongly suggest the existence of new physics beyond the Standard Model (BSM). To interpret such experimental results, numerous models have been proposed [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Among these, the low-energy supersymmetric model is one of the most widely discussed extensions of the SM [27, 28, 29]. Its primary motivation is to solve the gauge hierarchy problem in a natural way by introducing supersymmetric (SUSY) particles. These particles can also provide additional contributions to the muon g-2 anomaly, providing a potential solution to this puzzle. Moreover, assuming R-parity conservation, the lightest SUSY particle (LSP) is considered a compelling candidate for dark matter [30], capable of explaining the observed relic density in the universe.
@@
The Minimal Supersymmetric Standard Model (MSSM) [31, 32, 33, 34] is the simplest supersymmetric extension of the SM, known for its robust ability to explain various experimental anomalies [35, 36, 37, 38, 39, 40]. However, it includes over 100 free parameters, making their experimental determination extremely challenging. In contrast, the Constrained MSSM (CMSSM) [41, 42, 43, 44, 45, 46, 47] simplifies this by assuming unified values for the gaugino masses, scalar masses and trilinear coupling constants at the Grand Unified Theory (GUT) scale, represented as $M_{1/2}$, $M_{0}$ and $A_{0}$ respectively. This reduces the number of free parameters to five: $M_{0}$, $M_{1/2}$, $A_{0}$, $\tan\!\beta$ and the sign of $\mu$. The ratio $\tan\!\beta$ is defined as $V_{u}/V_{d}$, where $V_{u}$ and $V_{d}$ are the Higgs fields that give mass to the up-type quarks and the down-type quarks and leptons, respectively. This approach allows these five parameters to determine all low-energy soft SUSY breaking parameters through the renormalization group equations (RGEs).
The phenomenology of the CMSSM has been extensively studied [48, 49, 50]; however, by reducing the number of free parameters, the CMSSM offers a more constrained interpretation of the experimental data compared to the MSSM. In particular, it struggles to account for the anomalous magnetic moment of the muon while predicting a 125 GeV SM-like Higgs. This challenge arises because the masses of the light electroweakinos, higgsinos, sleptons, and neutrinos, which can contribute significantly to the muon anomalous magnetic moment, are predicted to be heavier than several hundred GeV in the CMSSM under these conditions.
The non-universal gaugino masses within the framework of SU(5) GUT offer a solution to the tensions observed between the CMSSM and current experimental data. Our study focuses on the scenario where $|M_{3}|\gg|M_{1}|,|M_{2}|$, referred to as $\tilde{g}$. We investigate the constraints imposed by current data on this parameter space, consider the annihilation mechanisms of dark matter relevant to this scenario, and explore the detection capabilities of future dark matter experiments. Finally, we evaluate the prospects of future dark matter direct detection experiments and collider studies of SUSY particles to probe this scenario.
@@
This paper is organized as follows. Section II provides a brief overview of the mechanism for generating non-universal gaugino masses in the $\tilde{g}$ framework within the context of the SU(5) GUT. Section III presents the numerical results obtained by applying the theoretical and experimental constraints, accompanied by a comprehensive theoretical interpretation. Finally, Section IV provides a summary of our conclusions.
II Non-universal gaugino masses in $\tilde{g}$
Based on different mechanisms of SUSY breaking, the CMSSM can be categorized into three main classes: Gravity Mediated SUSY Breaking (commonly referred to as minimal Supergravity or mSUGRA) [51, 52, 53, 54, 55], Anomaly Mediated SUSY Breaking (AMSB) [56, 57, 58], and Gauge Mediated SUSY Breaking (GMSB) [59, 60]. Among these, mSUGRA is a highly predictive framework that naturally accommodates a 125 GeV SM-like Higgs boson [61, 62]. In the mSUGRA scenario, the masses of the gauginos (bino, wino, and gluino) are unified at the GUT scale, expressed as $M_{1}=M_{2}=M_{3}=M_{1/2}$
The GUT of mSUGRA provides a comprehensive framework based on three arbitrary functions: the superpotential, the Kähler potential, and the gauge kinetic function. This framework allows for non-universalities in the soft SUSY-breaking parameters, including scalar masses, gaugino masses, and trilinear couplings [63, 64, 65, 66, 67, 68]. In this study, we focus on a scenario where the gaugino masses are non-universal, while other boundary conditions are maintained at the GUT scale.
The universal boundary conditions for soft parameters can be derived from specific assumptions about the Kähler potential and the gauge kinetic energy function [69]. Generally, the Kähler potential is assumed to be minimal, while the gauge kinetic terms for the vector supermultiplet can be expressed in a general form as:
$\mathcal{L}\supseteq\int d^{2}\theta\tau\left(W^{a}W^{a}+a_{1}\frac{S}{\Lambda}W^{a}W^{a}+b_{1}\frac{\Phi_{ab}}{\Lambda}W^{a}W^{b}\right),$ (5)
where $S$ and $\Phi_{ab}$ represent the singlet and non-singlet chiral superfields of the GUT group, respectively. The field $S$ can develop a vacuum expectation value (VEV) scaled by $\Lambda$ or lower, while non-singlet $\Phi_{ab}$ fields can generate F-term VEV. Soft SUSY-breaking gaugino masses are induced through these terms. Universal gaugino masses are generated by the $a_{1}$ term and associated with the $\langle F_{S}\rangle$, while non-universal gaugino masses are generated by the $b_{1}$ term and associated with the $\langle F_{\Phi_{ab}}\rangle$. To obtain non-universal gaugino masses, the non-singlet Higgs field $\Phi_{ab}$ must acquire non-vanishing F-term VEV. These operators can also contribute additional non-universalities to other soft SUSY breaking terms, such as trilinear couplings and sfermion masses.
Within the SU(5) GUT framework, the field $\Phi_{ab}$ can belong to any of the irreducible representations arising from the symmetric product of the adjoint representations. Specifically, the symmetric product of two adjoint representations, $\bm{24}\otimes\bm{24}$, can be decomposed into the following irreducible representations of SU(5):
$(\bm{24}\otimes\bm{24})_{\mathrm{symm}}=\bm{1}\oplus\bm{24}\oplus\bm{75}\oplus\bm{200}.$ (6)
Here, $\bm{24}$, $\bm{75}$, and $\bm{200}$ represent the chiral supermultiplets of SU(5). When $\Phi_{ab}$ is a singlet, as assumed in mSUGRA, the gaugino masses acquire a uniform value at the GUT scale. In contrast, if $\Phi_{ab}$ is a non-singlet belonging to a higher-dimensional representation of SU(5), the gaugino masses become non-uniform but remain interrelated through the representation invariants. For a representation denoted as $\bm{n}$, the three gaugino masses at the GUT scale possess distinct coefficients relative to the SUSY breaking mass parameter, and take the following form:
$M_{1,2,3}=C_{1,2,3}^{\bm{n}}\,M^{\bm{n}}_{1/2},$ (7)
where $M^{\bm{n}}_{1/2}$ is the soft gaugino mass arising from the F-term VEV in the representation $\bm{n}$. The coefficients $C^{\bm{n}}_{1,2,3}$ for specific representations are given as:
$C^{\bm{1}}_{1,2,3}$ $=(1,1,1),$ (8)
$C^{\bm{24}}_{1,2,3}$ $=(-1,-3,2),$
$C^{\bm{75}}_{1,2,3}$ $=(-5,3,1),$
$C^{\bm{200}}_{1,2,3}$ $=(10,2,1).$
For the Higgs fields belonging to the SU(5) GUT group with representations $\bm{24}$, $\bm{75}$, and $\bm{200}$, along with a gauge singlet $S$, the Lagrangian can be expressed as:
$\mathcal{L}=\int d^{2}\theta\left(\frac{\theta\tau}{4}W^{a}W^{b}\right)\frac{1}{\Lambda}\left[c_{0}S\delta_{ab}+c_{1}(\Phi^{\bm{24}})_{ab}+c_{2}(\Phi^{\bm{75}})_{ab}+c_{3}(\Phi^{\bm{200}})_{ab}\right].$ (9)
In particular, it is theoretically possible for the Higgs fields in these high-dimensional representations to acquire only the lowest component VEVs, while the singlet $S$ can acquire both its lowest component VEV and the F-term VEV. This configuration ensures the preservation of the $\mathrm{SU}(3)_{C}\times\mathrm{SU}(2)_{L}\times\mathrm{U}(1)_{Y}$ gauge symmetry, corresponding to the gauge boson superpartners, gluino ($\tilde{g}$), the wino ($\tilde{W}$), and the bino ($\tilde{B}$), respectively.
The VEVs of the Higgs fields $\Phi^{\bm{24}}$, $\Phi^{\bm{75}}$ and $\Phi^{\bm{200}}$ can be expressed in their respective adjoint representations, corresponding to $5\times 5$, $10\times 10$, and $15\times 15$ matrices. These matrices take on the following specific forms:
$\langle\Phi^{\bm{24}}\rangle=$ $\ v_{U}\sqrt{\frac{3}{5}}\mathrm{diag}\left(-\frac{1}{3},-\frac{1}{3},-\frac{1}{3},\frac{1}{2},\frac{1}{2}\right),$ (10)
$\langle\Phi^{\bm{75}}\rangle=$ $\ \dfrac{v_{U}}{2\sqrt{3}}\mathrm{diag}\left(\underbrace{1,\cdots,1}_{3},\underbrace{-1,\cdots,-1}_{6},3\right),$
$\langle\Phi^{\bm{200}}\rangle=$ $\ \dfrac{v_{U}}{2\sqrt{3}}\mathrm{diag}\left(\underbrace{1,\cdots,1}_{6},\underbrace{-2,\cdots,-2}_{6},\underbrace{2,\cdots,2}_{3}\right),$
where $v_{U}$ denotes the GUT breaking scale, which is assumed to have a universal value for all Higgs fields. In the most general scenarios, the singlet $S$ and the high-dimensional representations of Higgs fields are assumed to acquire universal F-term VEVs, $F_{\Lambda}$, under the assumption of trivial kinetic terms. Then $M_{1}$, $M_{2}$, and $M_{3}$ can get contributions from $S$, $\Phi^{\bm{24}}$, $\Phi^{\bm{75}}$ and $\Phi^{\bm{200}}$ at GUT scale and the ratio of gauginos is given by:
$M_{1}:M_{2}:M_{3}=$
$\left[c_{0}-\frac{c_{1}}{4\sqrt{15}}+\frac{5c_{2}}{4\sqrt{3}}+\frac{5c_{3}}{2\sqrt{3}}\right]:\left[c_{0}-\frac{3c_{1}}{4\sqrt{15}}-\frac{3c_{2}}{4\sqrt{3}}+\frac{c_{3}}{2\sqrt{3}}\right]:\left[c_{0}+\frac{c_{1}}{2\sqrt{15}}-\frac{c_{2}}{4\sqrt{3}}+\frac{c_{3}}{4\sqrt{3}}\right].$ (11)
@@ II Non-universal gaugino masses in $\tilde{g}$
Thus, an arbitrary ratio of $M_{1}:M_{2}:M_{3}$ at the GUT scale can be achieved by appropriately tuning the coefficients. A particularly interesting scenario arises when $M_{3}\gg M_{1},M_{2}$, ensuring that the radiative break of the electroweak symmetry is primarily driven by the gluino. This framework, often referred to as $\tilde{g}$-SUGRA [70], leads to rich phenomenological implications. In this case, the gluino significantly affects the RGEs of the squarks, contributing to their masses as the RGEs run from the GUT scale to the electroweak scale. Conversely, at the one-loop level, the gluino does not enter the RGEs of sleptons, leaving their masses unaffected. Such a scenario predicts heavy squarks, allowing them to escape current experimental constraints on squark masses. At the same time, the presence of relatively light sleptons can lead to a number of interesting phenomenological consequences.
III Results and discussion
@@
We use the SuSpect-2.52 [71] package to implement the theoretical and experimental constraints in our analysis, assuming $\text{sign}(\mu)=+1$. The remaining parameters are selected within the following ranges:
@@
  • The primary constraints on $M_{3}$, $A_{0}$, $M_{0}$ and $\tan\!\beta$ arise from theoretical considerations and Higgs data. In contrast, the primary constraints on $M_{1}$ and $M_{2}$ are imposed by dark matter data. The reasons for these differences will be discussed in more detail later.
@@
  • By combining the upper right and lower middle panels of Fig. 2, it appears that small values of $|A_{0}|$, $M_{3}$, and $M_{0}$ tend to provide a larger contribution to $\Delta a_{\mu}$. In particular, a sample with $M_{0}\sim 250\,\mathrm{GeV}$, $|A_{0}|\lesssim-1\,\mathrm{TeV}$, $M_{3}\sim 4\,\mathrm{TeV}$, and $\tan\!\beta\sim 10$ yields a sizable contribution to $\Delta a_{\mu}$ ($\sim 2.1\times 10^{-9}$). In SUSY, the primary contribution to $\Delta a_{\mu}$ comes from smuon interactions and is given by [88, 89]:
@@
  • From the upper left panel of Fig. 3, it is evident that the lightest neutralino mass $m_{\tilde{\chi}_{1}^{0}}$ is nearly proportional to $M_{2}$ in Class  A and Class  C, while it is nearly proportional to $M_{1}$ in Class  B. This behavior suggests the dominant component of the lightest neutralino $\tilde{\chi}_{1}^{0}$, as $M_{1}$ and $M_{2}$ are the mass parameters associated with the bino and wino, respectively. Furthermore, samples in Class  C predict a heavier $\tilde{\chi}_{1}^{0}$ compared to those in Class  A for the same $M_{2}$. This is attributed to the additional contribution of $M_{3}$ to $M_{2}$ at the soft SUSY scale.
@@
  • From the upper middle and right panels of Fig. 3, it can be observed that the lightest neutralino $\tilde{\chi}_{1}^{0}$ is predominantly wino-like in Class  A and Class  C, while it is predominantly bino-like in Class  B. Additionally, $\tilde{\chi}_{1}^{0}$ typically does not exhibit wino-bino mixing, except for samples near the lower edge of Class  A and Class  B.
@@
  • By combining the upper middle, upper right, and lower left panels of Fig. 3, it becomes clear that a larger dark matter relic density $\Omega h^{2}$ is only achieved when $\tilde{\chi}_{1}^{0}$ is predominantly bino-like. However, the $\Omega h^{2}$ predicted for bino-like $\tilde{\chi}_{1}^{0}$ is often excessively large and strongly constrained by experimental dark matter data. This limitation explains the scarcity of surviving samples in Class  B. In contrast, when $\tilde{\chi}_{1}^{0}$ exhibits wino-bino mixing, a significantly smaller $\Omega h^{2}$ can be achieved, although it remains larger than in scenarios where $\tilde{\chi}_{1}^{0}$ is predominantly wino-like.
@@
  • From the lower left and middle panels of Fig. 3, it can be observed that the mass of $\tilde{\chi}^{0}_{1}$ ($m_{\tilde{\chi}^{0}_{1}}$) and the mass of the lighter smuon $m_{\tilde{\tau}_{1}}$ are nearly degenerate when $\tilde{\chi}^{0}_{1}$ is bino-like. In contrast, $m_{\tilde{\tau}_{1}}$ becomes significantly larger than $m_{\tilde{\chi}^{0}_{1}}$ when $\tilde{\chi}^{0}_{1}$ is wino-like. This behavior provides insight into the annihilation mechanisms of $\tilde{\chi}^{0}_{1}$. When $\tilde{\chi}^{0}_{1}$ is bino-like, its annihilation is primarily mediated by interactions with $\tilde{\tau}_{1}$. Since $\tilde{\tau}_{1}$ is only slightly heavier than $\tilde{\chi}^{0}_{1}$, the annihilation process occurs relatively slowly, allowing for the accumulation of a sufficient relic density.
@@
  • From the lower left and right panels of Fig. 3, it can be observed that if $\tilde{\chi}_{1}^{0}$ is wino-like, its mass is nearly degenerate with the mass of the lightest chargino ($m_{\tilde{\chi}_{1}^{\pm}}$), which aligns with theoretical expectations. Consequently, $\tilde{\chi}^{0}_{1}$ can consistently co-annihilate with $\tilde{\chi}_{1}^{\pm}$ in wino-like scenarios. However, due to the high annihilation rate associated with this co-annihilation process, it is unable to produce a sufficient relic density ($\Omega h^{2}$). In contrast, if $\tilde{\chi}_{1}^{0}$ is bino-like, the mass of $\tilde{\chi}_{1}^{\pm}$ is significantly larger than that of $\tilde{\chi}_{1}^{0}$, making co-annihilation between the two particles inefficient.
@@
The dominant annihilation channels for each coannihilation scenario are as follows:
  • For $\tilde{\tau}\tilde{\chi}^{0}_{1}$ coannihilation, the primary channels are stau annihilation and neutralino-stau coannihilation.
  • For $\tilde{\chi}^{0}_{1}\tilde{\chi}^{\pm}_{1}$ coannihilation, the major channels are neutralino-chargino coannihilation and neutralino annihilation.
  • For $\tilde{\chi}^{\pm}_{1}\tilde{\chi}^{0}_{1}$ coannihilation, the main channels are chargino annihilation and neutralino-chargino coannihilation.
  • For $\tilde{\chi}^{0}_{1}$ annihilation, the dominant channel is neutralino annihilation.
@@
We provide detailed annihilation information for four benchmark points, including their top 10 annihilation channels and the relative contributions to $\sigma v$, in Table 1. Additional information about these benchmark points is summarized in Table 2.
@@
  • By comparing the upper middle and right panels of Fig.3, along with the left panel of Fig.5, it is evident that the primary annihilation mechanisms vary between classes:
    • In Class  A and Class  C, $\tilde{\chi}^{0}_{1}\tilde{\chi}^{\pm}_{1}$ coannihilation dominates, reflecting the tendency of wino-like $\tilde{\chi}^{0}_{1}$ to engage in this mechanism.
    • In Class  B, $\tilde{\tau}\tilde{\chi}^{0}_{1}$ coannihilation is the main mechanism, characteristic of bino-like $\tilde{\chi}^{0}_{1}$.
    • When $\tilde{\chi}^{0}_{1}$ exhibits bino-wino mixing, the dominant mechanism depends on the masses of $\tilde{\chi}^{0}_{1}$ and $\tilde{\chi}^{\pm}_{1}$, alternating between $\tilde{\chi}^{\pm}_{1}\tilde{\chi}^{0}_{1}$ coannihilation and $\tilde{\chi}^{0}_{1}$ annihilation.
  • From the middle and right panels of Fig. 5, it is observed that for samples with non-degenerate masses of $\tilde{\tau}$ and $\tilde{\chi}^{0}_{1}$, annihilation mechanisms include $\tilde{\chi}^{\pm}_{1}\tilde{\chi}^{0}_{1}$ coannihilation, $\tilde{\chi}^{0}_{1}\tilde{\chi}^{\pm}_{1}$ coannihilation, and $\tilde{\chi}^{0}_{1}$ annihilation. These mechanisms depend on the masses of $\tilde{\chi}^{0}_{1}$ and $\tilde{\chi}^{\pm}_{1}$, aligning with theoretical expectations.

### v2

@@ I Introduction
These phenomena are difficult to explain within the framework of the SM and strongly suggest the existence of new physics beyond the Standard Model (BSM). To interpret such experimental results, numerous models have been proposed [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Among these, low-energy supersymmetric models are most widely discussed extensions of the SM [25, 26, 27, 28].
However, direct searches for SUSY particles at the LHC [29, 30] have imposed stringent constraints on the gluino and squark masses, pushing them to the multi-TeV scale. This result challenges the original low-scale SUSY paradigm. Nevertheless, electroweakinos with masses at sub-TeV level remain viable, maintaining interest in supersymmetric models from both a phenomenological and experimental perspective.
@@
The Minimal Supersymmetric Standard Model (MSSM) [31, 32, 33, 34] is the simplest supersymmetric extension of the SM, known for its robust ability to explain various experimental anomalies [35, 36, 37, 38, 39, 40]. However, it includes over 100 free parameters, making their experimental determination extremely challenging. In contrast, assuming the unified values for the SUSY particle masses and coupling constants at the Grand Unified Theory (GUT) scale, the Constrained MSSM (CMSSM) [41, 42, 43, 44, 45, 46, 47] can reduces the free parameters to five.
The phenomenology of the CMSSM has been extensively studied [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70]. Among them, the gravity-mediated SUSY breaking (SUGRA) framework is highly predictive, which can naturally accommodates a 125 GeV SM-like Higgs boson [61, 62]. However, it struggles to account for the anomalous magnetic moment of the muon while predicting a 125 GeV SM-like Higgs. Then the Non-universal gaugino masses (NUGM) extension of CMSSM is proposed to ease the tension between CMSSM and the current experimental data, which allow $M_{1}\neq M_{2}\neq M_{3}$ [71, 72, 73, 74, 75]. In Ref. [72], a novel artificial neural network (ANN) method is used to scan the parameter space for studying the muon anomalous magnetic moment. Subsequently, the Monte Carlo method is used to investigate the capability of future colliders to cover the parameter space. Ref. [73] consider the scenarios involving non-universal gaugino masses, as well as non-universal slepton and Higgs masses, and then study the surviving status of benchmark points under the current experimental constraints from SUSY searches. In Ref. [75], a specialized mechanism based on the SU(5) Grand Unified Theory (GUT) is proposed to generate non-universal gaugino masses. The study provides a detailed discussion of dark matter constraints on this model. Ref. [74] analyzes the 2021 Fermilab experimental data on the muon anomalous magnetic moment, offering a comprehensive discussion of potential mechanisms to reconcile the tension between the muon anomalous magnetic moment and other experimental measurements.
In this work, we revisit the CMSSM with NUGM under current constraints and focus on the scenario with $|M_{3}|\gg|M_{1}|,|M_{2}|$. In such a scenario, heavy squarks and gluino can be predicted to escape the constraints on the direct search for SUSY particles. On the contrary, light electroweakinos are predicted to satisfy other experimental data. The constraints of muon anomalous magnetic moment, dark matter relic density, Higgs data, B physics, and SUSY particles direct search are considered. Additionally, we then investigate the constraints imposed by current data on this parameter space and study its surviving conditions. Then the annihilation mechanisms of dark matter are analyzed in detail.
Finally, we evaluate the prospects of future dark matter direct detection experiments and collider studies of SUSY particles to probe this scenario.
@@
This paper is organized as follows. Section II provides a brief overview of scanning parameter selection and the constraints considered. Section III presents the numerical results obtained by applying the theoretical and experimental constraints, accompanied by a comprehensive theoretical interpretation. Finally, Section IV provides a summary of our conclusions.
II Non-universal gaugino masses in $\tilde{g}$ SUGRA
The general CMSSM has trouble both in giving a 125 GeV Higgs and in dealing with the muon anomalous magnetic moment problem under current constraints. To address the longstanding tension between the CMSSM and experimental data, the NUGM extension of the CMSSM is proposed. One of the most interesting scenarios that has arisen is $M_{3}\gg M_{1},M_{2}$, which is also known as $\tilde{g}$-SUGRA.
@@ II Non-universal gaugino masses in $\tilde{g}$ SUGRA
In this case, the gluino can significantly affect the RGEs of the squarks, contributing to their masses as the RGEs run from the GUT scale to the electroweak scale. Conversely, at the one-loop level, the gluino does not enter into the RGEs of sleptons, leaving their masses unaffected. As a result, heavy squarks and gluino are predicted to evade the constraints on the direct search for SUSY particles, while light electroweakinos are predicted to satisfy other experimental data. In fact, the light smuon can provide a significant contribution to the g-2 to match the experimental result. Such a scenario predicts heavy squarks and gluino, allowing them to evade current experimental constraints on their masses. At the same time, the presence of relatively light sleptons can lead to a number of interesting phenomenological consequences.
@@
The scenario with $M_{3}\gg M_{1},M_{2}$ are considered in this work. We use the SuSpect-2.52 [76] package to implement the theoretical and experimental constraints in our analysis, assuming $\text{sign}(\mu)=+1$. The remaining parameters are selected within the following ranges:
@@ III Results and discussion
  • The primary constraints on $M_{3}$, $A_{0}$, $M_{0}$, and $\tan\!\beta$ arise from theoretical considerations and Higgs data. In contrast, the primary constraints on $M_{1}$ and $M_{2}$ are imposed by dark matter data. The reasons for these differences will be discussed in more detail later.
@@
Due to the use of more accurate calculation tools and the consideration of newer experimental constraints, the survival status of parameters in our study differs significantly from that in Ref. [75]. Specifically, samples with larger values of $\tan\beta$ ($\tan\beta\gtrsim 42$ or $\tan\beta\gtrsim 25$ with $M_{0}\approx 400\ \text{GeV}$) are excluded in our analysis, whereas they are allowed in Ref. [75].
@@
  • By combining the upper right and lower middle panels of Fig. 2, it appears that small values of $|A_{0}|$, $M_{3}$, and $M_{0}$ tend to provide a larger contribution to $\Delta a_{\mu}$. In particular, a sample with $M_{0}\sim 250\,\mathrm{GeV}$, $|A_{0}|\lesssim 1\,\mathrm{TeV}$, $M_{3}\sim 4\,\mathrm{TeV}$, and $\tan\!\beta\sim 10$ yields a sizable contribution to $\Delta a_{\mu}$ ($\sim 2.1\times 10^{-9}$). This result is consistent with Ref.[73], but differs from Ref.[72], as the latter considers a different set of theoretical assumptions and constraints. Our analysis incorporates additional constraints, leading to slight variations in the allowed parameter space. In SUSY, the primary contribution to $\Delta a_{\mu}$ comes from smuon interactions and is given by [91, 92]:
@@
  • From the upper left panel of Fig. 3, it is evident that the lightest neutralino mass $m_{\tilde{\chi}_{1}^{0}}$ is nearly proportional to $M_{2}$ for samples with $M_{1}/M_{2}\gtrsim 1.5$, while it is nearly proportional to $M_{1}$ for samples with $M_{1}/M_{2}\lesssim 1.5$. This behavior may suggests the dominant component of the lightest neutralino $\tilde{\chi}_{1}^{0}$, as $M_{1}$ and $M_{2}$ are the mass parameters associated with the bino and wino, respectively. Furthermore, samples with $M_{3}<0$ predict a heavier $\tilde{\chi}_{1}^{0}$ compared to samples with $M_{3}>0$ for the same $M_{2}$. This is attributed to the additional contribution of $M_{3}$ to $M_{2}$ at the soft SUSY scale.
@@
  • From the upper middle and right panels of Fig. 3, it can be observed that the lightest neutralino $\tilde{\chi}_{1}^{0}$ is predominantly wino-like when $M_{1}/M_{2}\gtrsim 1.5$, while it is predominantly bino-like when $M_{1}/M_{2}\lesssim 1.5$. Additionally, $\tilde{\chi}_{1}^{0}$ generally does not exhibit wino-bino mixing, except for samples with $M_{1}/M_{2}\approx 1.5$.
@@
  • By combining the upper middle, upper right, and lower left panels of Fig. 3, it becomes clear that a larger dark matter relic density $\Omega h^{2}$ is only achieved for samples with $M_{1}/M_{2}\lesssim 1.5$, where $\tilde{\chi}_{1}^{0}$ is predominantly bino-like or wino-bino mixing. However, the $\Omega h^{2}$ predicted for them is often excessively large and strongly constrained by experimental dark matter data. This limitation explains the scarcity of surviving samples with $M_{1}/M_{2}\lesssim 1.5$. In contrast, samples with $M_{1}/M_{2}\gtrsim 1.5$, where $\tilde{\chi}_{1}^{0}$ is predominantly wino-like, tend to predict a small $\Omega h^{2}$ to escape the dark matter relic density data.
@@
  • From the lower left and middle panels of Fig. 3, it can be observed that the mass of the lightest neutralino $m_{\tilde{\chi}^{0}_{1}}$ and the mass of the lighter smuon $m_{\tilde{\tau}_{1}}$ are nearly degenerate when $M_{1}/M_{2}\lesssim 1.5$. In contrast, $m_{\tilde{\tau}_{1}}$ becomes significantly larger than $m_{\tilde{\chi}^{0}_{1}}$ when $M_{1}/M_{2}\gtrsim 1.5$. This behavior may provide insight into the annihilation mechanisms of $\tilde{\chi}^{0}_{1}$. When $\tilde{\chi}^{0}_{1}$ is bino-like, its annihilation may be primarily mediated by interactions with $\tilde{\tau}_{1}$. Since $\tilde{\tau}_{1}$ is only slightly heavier than $\tilde{\chi}^{0}_{1}$, the annihilation process occurs relatively slowly, allowing for the accumulation of a sufficient relic density.
@@
  • From the lower left and right panels of Fig. 3, it can be observed that when $M_{1}/M_{2}\gtrsim 1.5$, the mass of the lightest neutralino $m_{\tilde{\chi}^{0}_{1}}$ nearly degenerate with the mass of the lightest chargino $m_{\tilde{\chi}_{1}^{\pm}}$, which aligns with theoretical expectations for wino-like samples. Consequently, $\tilde{\chi}^{0}_{1}$ can consistently co-annihilate with $\tilde{\chi}_{1}^{\pm}$ for these samples. However, due to the high annihilation rate associated with this co-annihilation process, it fails to produce a sufficient relic density ($\Omega h^{2}$). In contrast, when $\tilde{\chi}_{1}^{0}$ is bino-like, the mass of $\tilde{\chi}_{1}^{\pm}$ is significantly larger than that of $\tilde{\chi}_{1}^{0}$, making co-annihilation between the two particles inefficient.
@@
The dominant annihilation channels for these scenarios are as follows:
  • $\tilde{\tau}\tilde{\chi}^{0}_{1}$ coannihilation: the primary channels are stau annihilation and neutralino-stau coannihilation.
  • $\tilde{\chi}^{0}_{1}\tilde{\chi}^{\pm}_{1}$ coannihilation: the major channels are neutralino-chargino coannihilation and neutralino annihilation.
  • $\tilde{\chi}^{\pm}_{1}\tilde{\chi}^{0}_{1}$ coannihilation: the main channels are chargino annihilation and neutralino-chargino coannihilation.
  • $\tilde{\chi}^{0}_{1}$ annihilation: the dominant channel is neutralino annihilation.
@@
We provide detailed annihilation information for four benchmark points, including their top 10 annihilation channels and the relative contributions to $\sigma v$, in Table 1. Additional information about these benchmark points is summarized in Table 2. In contrast to Refs. [72, 73, 74], which focus primarily on parameter space constraints, our study includes a comprehensive analysis of the dark matter annihilation mechanisms, providing a more detailed understanding of their phenomenological implications.
@@
  • By comparing the upper middle and right panels of Fig. 3, along with the left panel of Fig. 5, one can see that the dominant annihilation mechanisms of the samples with $M_{1}/M_{2}\gtrsim 1.5$ are $\tilde{\chi}^{0}_{1}\tilde{\chi}^{\pm}_{1}$ coannihilation, which reflects the tendency of wino-like $\tilde{\chi}^{0}_{1}$ to engage in this mechanism. In contrast, for the bino-like samples with $M_{1}/M_{2}\lesssim 1.5$, $\tilde{\tau}\tilde{\chi}^{0}_{1}$ coannihilation is the main mechanism. The dominant mechanism can also be $\tilde{\chi}^{\pm}_{1}\tilde{\chi}^{0}_{1}$ coannihilation and $\tilde{\chi}^{0}_{1}$ annihilation when $M_{1}/M_{2}\approx 1.5$ or $M_{1}\approx 1.5M_{2}+200\ \mathrm{GeV}$, depending on the relative masses $\tilde{\chi}^{0}_{1}$ and $\tilde{\chi}^{\pm}_{1}$.
  • From the middle and right panels of Fig. 5, it is observed that for samples with degenerate masses of $\tilde{\tau}$ and $\tilde{\chi}^{0}_{1}$, the main annihilation mechanism is $\tilde{\tau}\tilde{\chi}^{0}_{1}$ coannihilation. In contrast, for samples with non-degenerate masses of $\tilde{\tau}$ and $\tilde{\chi}^{0}_{1}$, annihilation mechanisms include $\tilde{\chi}^{\pm}_{1}\tilde{\chi}^{0}_{1}$ coannihilation, $\tilde{\chi}^{0}_{1}\tilde{\chi}^{\pm}_{1}$ coannihilation, and $\tilde{\chi}^{0}_{1}$ annihilation. The particular annihilation mechanism depends on the mass relation between $\tilde{\chi}^{0}_{1}$ and $\tilde{\chi}^{\pm}_{1}$. Specifically, the main annihilation mechanism is $\tilde{\chi}^{\pm}_{1}\tilde{\chi}^{0}_{1}$ coannihilation or $\tilde{\chi}^{0}_{1}$ annihilation when $m_{\tilde{\chi}^{\pm}_{1}}/m_{\tilde{\chi}^{0}_{1}}\approx 1$ or $m_{\tilde{\chi}^{\pm}_{1}}/m_{\tilde{\chi}^{0}_{1}}\approx 1.1$. And the main annihilation mechanisms can be $\tilde{\chi}^{0}_{1}\tilde{\chi}^{\pm}_{1}$ coannihilation when $1\lesssim m_{\tilde{\chi}^{\pm}_{1}}/m_{\tilde{\chi}^{0}_{1}}\lesssim 1.1$.